# Magnon transmission across $\nu=1|-1|1$ mono-layer graphene junction as a probe of electronic structure

### v1

@@ Abstract
We study magnon transmission across gate-controlled junctions in the $n=0$ manifold of Landau levels in monolayer graphene, allowing for both spin and valley Zeeman fields. We show that by tuning the external perpendicular magnetic field, the phase in the intermediate region of a $11|\nu_{m}|$ 1 sandwich geometry can be changed, due to which the magnon transmission can be switched between fully transmitting and fully blocked. Our analysis, along with the experimental measurements, can be used to determine the anisotropic couplings in the sample.
@@ I Introduction
In addition to the long-range Coulomb interaction, there are lattice scale interactions that break the $SU(4)$ symmetry and compete with the Zeeman and valley Zeeman couplings in selecting the ground state [35, 36, 37, 38, 33, 17]. It has long been known that at the four-fermion level, the low-energy effective theory must conserve the number of electrons in each valley separately, leading to a $U(1)_{v}$ symmetry [35]. In a seminal work, Kharitonov proposed an ultra-short-range (USR) model [39, 40] for the residual interactions which has two couplings, a valley Ising coupling $u_{z}$, and a valley $XY$ coupling $u_{\perp}$, both expected to scale linearly with $B_{\perp}$. He then solved this model in the Hartree-Fock (HF) approximation to obtain a phase diagram in the $u_{z}-u_{\perp}$ parameter space. At charge neutrality, the following quantum Hall phases are seen [39]: A fully spin-polarized quantum spin Hall insulator (FM), a canted antiferromagnet (CAF), a bond-ordered (BO) state, presumably with a Kekulé distortion (and thus often called KD, though we will use the notation BO), and a charge density wave (CDW). It is believed that at a purely perpendicular field, the system is either in a BO state or a CAF state, and upon increasing $E_{Z}$ it undergoes a transition to the fully spin-polarized state. More recently, even more complex phase diagrams have been proposed both for $\nu=0$ [41, 42, 43, 44] and $\nu=\pm 1$ [45, 46] systems by relaxing the ultra-short-range assumption for the anisotropic residual couplings.
@@
Many experimental techniques have been used to probe the $\nu=0$ state. Apart from transport [22, 23, 24, 25], the state has been probed by scanning tunneling microscopy/spectroscopy (STM) [30, 31, 32] and by the transmission of magnons [26, 27, 28, 29]. Magnons are collective excitations which carry spin, and are always present in systems which have a nonzero spin polarization. These two techniques probe different order parameters. While magnon transmission probes whether the state in question has any spin polarization, current STM experiments are not spin-resolved. They detect charge density at the atomic scale, and can thus detect the presence of charge and/or bond order. STM experiments ubiquitously show bond order and CDW order [30, 31, 32]. As an aside, we note that while bond order and CAF order do not coexist in Kharitonov’s phase diagram [39], removing the restriction of ultra-short range interactions allows them to coexist [41, 42, 43, 44] at $\nu=0$. Based on a combination of STM and magnon transmission experiments, the current consensus is that at low $B_{\perp}$, the system is a spin singlet and has bond order. As the field $B_{\perp}$ increases, the couplings $u_{z}$ and $u_{\perp}$ increase, while $E_{V}$ remains the same. As detected by magnon transmission [29], the system makes a transition into a magnetic state, presumably the CAF phase, at some critical value of $B_{\perp}$. Based on Kharitonov’s phase diagram this puts physical systems in the region of the parameter space where $u_{z}>0$ while $u_{\perp}<0$.
@@
Our goal in this work is to thoroughly examine a much simpler system, which is a “sandwich” of $\nu=1$, $\nu=-1$ and $\nu=1$, denoted as $1|-1|1$. Such a system has been examined before in the limiting case when $E_{V}=0$ [47]. We study it in full generality in the neighborhood of the physical region of the $u_{z},\ u_{\perp}$ parameter space with nonzero $E_{Z},\ E_{V}$. Real samples of graphene are believed to have $u_{z}>0,\ u_{\perp}<0$. We will examine both $u_{z}>|u_{\perp}|$, which puts the system at $E_{Z}=E_{V}=0$ in the AF phase, and $u_{z}<|u_{\perp}|$, for which the system is in the bond-ordered phase at $E_{Z}=E_{V}=0$. Since $E_{Z},u_{z},u_{\perp}$ all scale linearly with $B_{\perp}$ while $E_{V}$ remains constant, one can access many different regimes simply by varying $B_{\perp}$. The first step is to examine the Hartree-Fock (HF) ground states for each of $\nu=\pm 1$. As we will show, the ordering of the HF energies in the $\nu=\pm 1$ states, which depends on the coupling constants $u_{z},\ u_{\perp}$ as well as $E_{Z},\ E_{V}$, plays a key role in the nature of the interface between $\nu=1$ and $\nu=-1$. The middle layer of the sandwich can undergo transitions between different ground states as $B_{\perp}$ is varied. Note that magnon scattering in a skyrmion crystal with a $11|1\pm\delta\nu|$ 1 sandwich also has been studied earlier [48].
@@
Experimentally, magnons can be generated at the the contacts in a $\nu=1$ system [26, 49, 28, 27, 29, 50] by creating a potential difference between the co-propagating edge channels of opposite spins at an edge between $\nu=2$ and $\nu=1$. When the bias voltage between the two channels exceeds the spin flip energy $2E_{Z}$, a magnon is created by a spin-flip process, which can then be transmitted through the $11|\nu_{m}|$ 1 junction, where $\nu_{m}$ is the filling fraction in the middle region. The magnon transmission probability through the sandwich is studied as a function of the incident magnon energy, either by local conductance measurements or non-local voltage measurements. In an earlier work [26] (believed to be for $E_{V}\approx 0$) it was found that for both configurations $1|1|1$ and $1|-1|1$, the magnons were largely transmitted, while for $1|0|1$, they were largely reflected for energies near threshold ($2E_{Z}$). Although this can be understood through kinematic constraints  [47], later experiments [28] show magnons are largely reflected for $1|-1|1$. More recent experiments [29] show that the transmission across the junction can be changed by tuning the external perpendicular magnetic field $B_{\perp}$. Thus, even for the simpler system, the experimental situation is far from clear. This is a strong motivation for our detailed study including all allowed parameters. Our main result, in this paper, is to show that magnon transmission through the $1|-1|1$ ‘sandwich’ can detect various phases in the intermediate region and the phase transitions, as a function of $B_{\perp}$.
@@
Another important motivation for us is the possibility of determining all the coupling parameters using magnon transmission across the sandwich. The idea is as follows: The value of $E_{V}$ can be determined by zero-$B_{\perp}$ measurements of the gap [18]. Using this and the known value of $E_{Z}$, one is left with the free parameters $u_{z},\ u_{\perp}$. The three parameters $E_{Z},u_{z},u_{\perp}$ all scale linearly with $B_{\perp}$. The Coulomb interaction scales with $\sqrt{B_{\perp}}$, whereas $E_{V}$ is independent of $B_{\perp}$. Thus, by varying $B_{\perp}$ one may be able to vary the ratios of these parameters over a wide range. The ratios determine the ordering of the one-body levels, and consequently the structure of the junction, and the middle region of the sandwich. Of particular interest are specific ratios of parameters where phase transitions in the order parameters of the middle region occur, which are reflected in the magnon transmission probability. Thus, magnon transmission over a wide range of $B_{\perp}$ would ideally allow us to determine $u_{z}/B_{\perp}$ and $u_{\perp}/B_{\perp}$ for a given sample.
@@ II The bulk Hamiltonian and Hartree Fock ground state
where $\phi(k,\vec{q})=\ell^{2}(-q_{x}k+\frac{1}{2}q_{x}q_{y})$ and $\vec{c}_{k}=(c_{kK{\uparrow}},c_{kK{\downarrow}},c_{kK^{\prime}{\uparrow}},c_{kK^{\prime}{\downarrow}})^{T}$. The matrices $\tau_{\alpha}$ and $\sigma_{\alpha}$, $\alpha=0,x,y,z$, with 00 denoting the identity matrix, are Pauli matrices acting in the valley and spin spaces respectively. More explicitly, $\tau_{\alpha}\equiv\tau_{\alpha}\otimes\sigma_{0}$. Here, $V(q_{x},q_{y})=\frac{E_{c}}{\sqrt{q^{2}\ell^{2}+q_{0}^{2}}}$ is the screened Coulomb interaction. We have used the ultra-short-range (USR) assumption for the residual anisotropic interactions [39], implying that $v_{\alpha}(q)$ are independent of momentum; $v_{\alpha}(q)=2\pi\ell^{2}u_{\alpha}$. The valley $XY$ coupling is given by $u_{x}=u_{y}=u_{\perp}$. As seen in earlier work [41, 42, 43, 44], relaxing the USR assumption does lead to the appearance of new phases at $\nu=0$. However, it does not seem to have a strong effect on the phases of $\nu=\pm 1$ in the physical region of the parameter space [45, 51, 46], which is why we continue to use the USR assumption here. We will keep the Coulomb screening wavevector $q_{0}$ small in our analysis. $E_{Z}$ is the spin Zeeman term, which denotes the coupling of the electron spin with the external magnetic field, $E_{Z}\propto\mu_{b}{B_{\perp}}$, where $\mu_{b}$ is the Bohr magneton. $E_{V}$ is the valley Zeeman/sublattice potential term, which breaks the valley degeneracy in MLG and favors the $K$ valley over the $K^{\prime}$ valley in the noninteracting limit. As explained in the introduction, this term is usually generated from the partial misalignment of the substrate layer, (such as hexagonal boron nitrate (hBN) layer) with the graphene layer [18, 19, 20, 21].
@@
$E_{K^{\prime}\uparrow}=-V_{ex}+E_{V}-E_{Z}-2u_{z}-2u_{\perp},$
$E_{K^{\prime}\downarrow}=E_{V}+E_{Z}-u_{z}-2u_{\perp}$ (4)
@@
$E_{K\downarrow}-E_{K^{\prime}\uparrow}$ $=$ $2({B_{\perp}}[E_{Z}^{0}+(u_{z}^{0}-|u_{\perp}^{0}|)]-E_{V})$
$E_{K\uparrow}-E_{K\downarrow}$ $=$ $2{B_{\perp}}(|u_{\perp}^{0}|-E_{Z}^{0})$ (6)
@@
Since there are multiple energy scales involved, we will consider two illustrative cases, leaving the detailed investigation of all the possibilities to later sections. In both cases we will look only at what is believed to be the physical region of anisotropic couplings, given by $u_{z}>0,\ u_{\perp}<0$. In Fig.  (1) we focus on the case $u_{z}>|u_{\perp}|$. Assuming that the field is perpendicular, and that $u_{z}^{0}-|u_{\perp}^{0}|>E_{z}^{0}$. Finally, assuming a partially aligned HBN substrate, we take $E_{V}>u_{z}^{0}-|u_{\perp}^{0}|$. As seen in Fig.  (1), there are three different orderings as a function of $B_{\perp}$. For $B_{\perp}<E_{V}/u_{z}^{0}$ (roughly $1.3T$ for the parameters chosen) the ordering of the levels is $E_{{K\downarrow}}<E_{{K\uparrow}}<E_{{K^{\prime}\uparrow}}$. For intermediate values $\frac{E_{V}}{E_{Z}^{0}+(u_{z}^{0}-|u_{\perp}^{0}|)}>B_{\perp}>\frac{E_{V}}{u_{z}^{0}}$, the ordering becomes $E_{{K\downarrow}}<E_{{K^{\prime}\uparrow}}<E_{{K\uparrow}}$. Finally, for large values of $B_{\perp}$, beyond roughly $3.4T$ for the parameters chosen, the ordering is $E_{{K^{\prime}\uparrow}}<E_{{K\downarrow}}<E_{{K\uparrow}}$.
@@
Figure 1: HF energies of the occupied levels for bulk $\nu=1$ as a function of ${B_{\perp}}/{B_{\perp}}^{0}$ with $V_{ex}=0$$E_{Z}^{0}=0.5$$E_{V}=5$$u_{z}^{0}=4$ and $u_{\perp}^{0}=-3$ i.e. the case when $E_{V}>~{}(u_{z}^{0}-|u_{\perp}^{0}|)>~{}E_{Z}^{0}$. For ${B_{\perp}}/{B_{\perp}}^{0}\simeq 1$, the ordering of the occupied levels is given by $E_{K\downarrow}<E_{K\uparrow}<E_{K^{\prime}\uparrow}$. As we increase ${B_{\perp}}$, the energy of the state ${K^{\prime}\uparrow}$ decreases linearly while the other energies increase, thus changing the order of the occupied levels.
@@
As a second illustrative example, let us consider a case where, at $E_{Z}=E_{V}=0$, the system would have been in the BO state in the physical region of anisotropic couplings, that is, $u_{z}<|u_{\perp}|$. We will consider a sceniario similar to the one in the previous paragraph, with $E_{V}>(|u_{\perp}^{0}|-u_{z}^{0})>E_{Z}^{0}$. Now there are only two different types of orderings of single-particle energy levels, as seen in Fig.  (2). For $B_{\perp}<E_{V}/u_{z}^{0}$ (roughly 2.5 for the parameters chosen) the ordering is $E_{k\downarrow}<E_{K\uparrow}<E_{K^{\prime}\uparrow}$. For larger $B_{\perp}$ the ordering switches to $E_{k\downarrow}<E_{K^{\prime}\uparrow}<E_{K\uparrow}$.
@@
Figure 2: HF energies of the occupied levels for bulk $\nu=1$ as a function of ${B_{\perp}}/{B_{\perp}}^{0}$ with $V_{ex}=0$$E_{Z}^{0}=0.5$$E_{V}=5$$u_{z}^{0}=2$ and $u_{\perp}^{0}=-3$ i.e. the case when $E_{V}>(|u_{\perp}^{0}|-u_{z}^{0})>E_{Z}^{0}$. For ${B_{\perp}}/{B_{\perp}}^{0}\simeq 1$, the ordering of the occupied levels are again $E_{K\downarrow}<E_{K\uparrow}<E_{K^{\prime}\uparrow}$, but as we increase ${B_{\perp}}$, the energies of all the occupied levels increase. We find that the ordering of the occupied levels changes only once, in contrast to Fig. (1) where it changes twice.
@@
From these examples it is clear that for fixed $E_{V}$ and purely perpendicular field, the ratio of $u_{z}$ and $u_{\perp}$ (that is, whether $\frac{u_{z}}{|u_{\perp}|}>1$ or $\frac{u_{z}}{|u_{\perp}|}<1$) as well as the magnitude of $B_{\perp}$ control the energetic ordering of filled levels at $\nu=1$. While any static ground state average will not be affected by the ordering of the filled levels, we will show that this ordering has drastic effects on the junctions, and thence on the magnon transmission.

### v2

@@ Abstract
We study magnon transmission across gate-controlled junctions in the $n=0$ manifold of Landau levels in monolayer graphene, in the presence of both spin and valley Zeeman fields. Specifically, we consider the $1|-1|1$ sandwich geometry. The nature of the interfaces between regions of different filling turns out to be crucial for magnon transmission. Using the Hartree-Fock approximation, we find that either the spin or the valley degrees of freedom of the occupied one-body states rotate across the interfaces. If the interfaces exhibit spin rotation, magnon transmission is suppressed at high energies, while if the interfaces have valley rotation, magnon transmission becomes perfect at high energies. The valley Zeeman coupling, which arises from partial alignment with the encapsulating Boron Nitride, is independent of perpendicular magnetic field $B$, while the spin Zeeman and other anisotropic couplings scale linearly with $B$. This allows the tuning of the relative strength of the valley Zeeman coupling in situ by varying $B$, which can drive phase transitions of the interfaces between spin-rotated and valley-rotated phases, leading to magnon transmission being either vanishing or perfect at high energies. Our analysis, along with the experimental measurements, can be used to determine the anisotropic couplings in the sample.
@@ I Introduction
In addition to the long-range Coulomb interaction, there are lattice scale interactions that break the $SU(4)$ symmetry and compete with the Zeeman and valley Zeeman couplings in selecting the ground state [35, 36, 37, 38, 33, 17]. It has long been known that at the four-fermion level, the low-energy effective theory must conserve the number of electrons in each valley separately, leading to a $U(1)_{v}$ symmetry [35]. In a seminal work, Kharitonov proposed an ultra-short-range (USR) model [39, 40] for the residual interactions which has two couplings, a valley Ising coupling $u_{z}$, and a valley $XY$ coupling $u_{xy}$, both expected to scale linearly with $B_{\perp}$. He then solved this model in the Hartree-Fock (HF) approximation to obtain a phase diagram in the $u_{z}-u_{xy}$ parameter space. At charge neutrality, the following quantum Hall phases are seen [39]: A fully spin-polarized quantum spin Hall insulator (FM), a canted antiferromagnet (CAF), a bond-ordered (BO) state, presumably with a Kekulé distortion (and thus often called KD, though we will use the notation BO), and a charge density wave (CDW). It is believed that at a purely perpendicular field, the system is either in a BO state or a CAF state, and upon increasing $E_{Z}$ it undergoes a transition to the fully spin-polarized state. More recently, even more complex phase diagrams have been proposed both for $\nu=0$ [41, 42, 43, 44] and $\nu=\pm 1$ [45, 46] systems by relaxing the ultra-short-range assumption for the anisotropic residual couplings.
@@
Many experimental techniques have been used to probe the $\nu=0$ state. Apart from transport [22, 23, 24, 25], the state has been probed by scanning tunneling microscopy/spectroscopy (STM) [30, 31, 32] and by the transmission of magnons [26, 27, 28, 29]. Magnons are collective excitations which carry spin, and are always present in systems which have a nonzero spin polarization. These two techniques probe different order parameters. While magnon transmission probes whether the state in question has any spin polarization, current STM experiments are not spin-resolved. They detect charge density at the atomic scale, and can thus detect the presence of charge and/or bond order. STM experiments ubiquitously show bond order and CDW order [30, 31, 32]. As an aside, we note that while bond order and CAF order do not coexist in Kharitonov’s phase diagram [39], removing the restriction of ultra-short range interactions allows them to coexist [41, 42, 43, 44] at $\nu=0$. Based on a combination of STM and magnon transmission experiments, the current consensus is that at low $B_{\perp}$, the system is a spin singlet and has bond order. As the field $B_{\perp}$ increases, the couplings $u_{z}$ and $u_{xy}$ increase, while $E_{V}$ remains the same. As detected by magnon transmission [29], the system makes a transition into a magnetic state, presumably the CAF phase, at some critical value of $B_{\perp}$. Based on Kharitonov’s phase diagram this puts physical systems in the region of the parameter space where $u_{z}>0$ while $u_{xy}<0$.
@@
Our goal in this work is to thoroughly examine a much simpler system, which is a “sandwich” of $\nu=1$, $\nu=-1$ and $\nu=1$, denoted as $1|-1|1$. Such a system has been examined before in the limiting case when $E_{V}=0$ [47]. We study it in full generality in the neighborhood of the physical region of the $u_{z},\ u_{xy}$ parameter space with nonzero $E_{Z},\ E_{V}$. Real samples of graphene are believed to have $u_{z}>0,\ u_{xy}<0$. We will examine both $u_{z}>|u_{xy}|$, which puts the system at $E_{Z}=E_{V}=0$ in the AF phase at $\nu=0$, and $u_{z}<|u_{xy}|$, for which the system is in the bond-ordered phase at $E_{Z}=E_{V}=0$ at $\nu=0$. Since $E_{Z},u_{z},u_{xy}$ all scale linearly with $B_{\perp}$ while $E_{V}$ remains constant, one can access many different regimes simply by varying $B_{\perp}$. The first step is to examine the Hartree-Fock (HF) ground states for each of $\nu=\pm 1$. It is important to note that for $\nu=\pm 1$ ultra-short-range interactions are unable to pick out a unique bulk ground state if $E_{Z}=E_{V}=0$. We will always have both types of Zeeman fields nonzero in what follows, because in this situation, the bulk ground states at $\nu=\pm 1$ are unique. As we will show, the ordering of the HF energies in the $\nu=\pm 1$ states, which depends on the coupling constants $u_{z},\ u_{xy}$ as well as $E_{Z},\ E_{V}$, plays a key role in the nature of the interface between $\nu=1$ and $\nu=-1$. The middle layer of the sandwich can undergo transitions between different ground states as $B_{\perp}$ is varied. Note that magnon scattering in a skyrmion crystal with a $11|1\pm\delta\nu|$ 1 sandwich also has been studied earlier [48].
@@
Experimentally, magnons can be generated at the the contacts in a $\nu=1$ system [26, 49, 28, 27, 29, 50] by creating a potential difference between the co-propagating edge channels of opposite spins at an edge between $\nu=2$ and $\nu=1$. When the bias voltage between the two channels exceeds the spin flip energy $2E_{Z}$, a magnon is created by a spin-flip process, which can then be transmitted through the $11|\nu_{m}|$ 1 junction, where $\nu_{m}$ is the filling fraction in the middle region. The magnon transmission probability through the sandwich is studied as a function of the incident magnon energy, either by local conductance measurements or non-local voltage measurements. In an earlier work [26] (believed to be for $E_{V}\approx 0$) it was found that for both configurations $1|1|1$ and $1|-1|1$, the magnons were largely transmitted, while for $1|0|1$, they were largely reflected for energies near threshold ($2E_{Z}$). Although this can be understood through kinematic constraints  [47], later experiments [28] show magnons are largely reflected for $1|-1|1$, perhaps because of oblique incidence at the interface. More recent experiments [29] performed in the range $10<\nu<$ 1 of the middle region show that the transmission across the junction can be changed by tuning the external perpendicular magnetic field $B_{\perp}$. Thus, even for the simpler system, the experimental situation is far from clear. This is a strong motivation for our detailed study including all allowed parameters. Our main result in this paper is to show that magnon transmission through the $1|-1|1$ sandwich can detect various phases in the intermediate region as a function of $B_{\perp}$.
@@
Another important motivation for us is the possibility of determining all the coupling parameters using magnon transmission across the sandwich. The idea is as follows: The value of $E_{V}$ can be determined by zero-$B_{\perp}$ measurements of the gap [18]. Using this and the known value of $E_{Z}$, one is left with the free parameters $u_{z},\ u_{xy}$. The three parameters $E_{Z},u_{z},u_{xy}$ all scale linearly with $B_{\perp}$. The Coulomb interaction scales with $\sqrt{B_{\perp}}$, whereas $E_{V}$ is independent of $B_{\perp}$. Thus, by varying $B_{\perp}$ one is able to vary the ratios of these parameters over a wide range. The ratios determine the ordering of the one-body levels, and consequently the structure of the junction, and the order parameters in the middle region of the sandwich. Of particular interest are specific ratios of the coupling constants where phase transitions in the order parameters of the middle region occur, which are reflected in the magnon transmission probability. Thus, magnon transmission over a wide range of $B_{\perp}$ would ideally allow us to determine $u_{z}/B_{\perp}$ and $u_{xy}/B_{\perp}$ for a given sample.
@@ II The bulk Hamiltonian and Hartree Fock ground state
where $\phi(k,\vec{q})=\ell^{2}(-q_{x}k+\frac{1}{2}q_{x}q_{y})$ and $\vec{c}_{k}=(c_{kK{\uparrow}},c_{kK{\downarrow}},c_{kK^{\prime}{\uparrow}},c_{kK^{\prime}{\downarrow}})^{T}$. The matrices $\tau_{\alpha}$ and $\sigma_{\alpha}$, $\alpha=0,x,y,z$, with 00 denoting the identity matrix, are Pauli matrices acting in the valley and spin spaces respectively. More explicitly, $\tau_{\alpha}\equiv\tau_{\alpha}\otimes\sigma_{0}$. Here, $V(q_{x},q_{y})=\frac{E_{c}}{\sqrt{q^{2}\ell^{2}+q_{0}^{2}}}$ is the screened Coulomb interaction. We have used the ultra-short-range (USR) assumption for the residual anisotropic interactions [39], implying that $v_{\alpha}(q)$ are independent of momentum; $v_{\alpha}(q)=2\pi\ell^{2}u_{\alpha}$. The valley $XY$ coupling is given by $u_{x}=u_{y}=u_{xy}$. As seen in earlier work [41, 42, 43, 44], relaxing the USR assumption does lead to the appearance of new phases at $\nu=0$. However, as long as $E_{Z},E_{V}>0$, which we will assume throughout this work, the USR assumption does not seem to have a strong effect on the phases of $\nu=\pm 1$ in the physical region of the parameter space [45, 51, 46], which is why we continue to use the USR assumption here. We will keep the Coulomb screening wavevector $q_{0}$ small in our analysis. $E_{Z}$ is the spin Zeeman term, which denotes the coupling of the electron spin with the external magnetic field, $E_{Z}\propto\mu_{b}{B_{\perp}}$, where $\mu_{b}$ is the Bohr magneton. $E_{V}$ is the valley Zeeman/sublattice potential term, which breaks the valley degeneracy in MLG and favors the $K$ valley over the $K^{\prime}$ valley in the noninteracting limit. As explained in the introduction, this term is usually generated from the partial misalignment of the substrate layer, (such as HBN layer) with the graphene layer [18, 19, 20, 21].
@@
$E_{K^{\prime}\uparrow}=-V_{ex}+E_{V}-E_{Z}-2u_{z}-2u_{xy},$
$E_{K^{\prime}\downarrow}=E_{V}+E_{Z}-u_{z}-2u_{xy}$ (4)
@@
$E_{K\downarrow}-E_{K^{\prime}\uparrow}$ $=$ $2({B_{\perp}}[E_{Z}^{0}+(u_{z}^{0}-|u_{xy}^{0}|)]-E_{V})$
$E_{K\uparrow}-E_{K\downarrow}$ $=$ $2{B_{\perp}}(|u_{xy}^{0}|-E_{Z}^{0})$ (6)
@@
Since there are multiple energy scales involved, we will consider two illustrative cases, leaving the detailed investigation of all the possibilities to later sections. In both cases we will look only at what is believed to be the physical region of anisotropic couplings, given by $u_{z}>0,\ u_{xy}<0$. In Fig.  (1) we focus on the case $u_{z}>|u_{xy}|$. Assuming that the field is perpendicular, and that $u_{z}^{0}-|u_{xy}^{0}|>E_{Z}^{0}$. Finally, assuming a partially aligned HBN substrate, we take $E_{V}>u_{z}^{0}-|u_{xy}^{0}|$. As seen in Fig.  (1), there are three different orderings as a function of $B_{\perp}$. For $B_{\perp}<E_{V}/u_{z}^{0}$ (roughly $1.3T$ for the parameters chosen) the ordering of the levels is $E_{{K\downarrow}}<E_{{K\uparrow}}<E_{{K^{\prime}\uparrow}}$. For intermediate values $\frac{E_{V}}{E_{Z}^{0}+(u_{z}^{0}-|u_{xy}^{0}|)}>B_{\perp}>\frac{E_{V}}{u_{z}^{0}}$, the ordering becomes $E_{{K\downarrow}}<E_{{K^{\prime}\uparrow}}<E_{{K\uparrow}}$. Finally, for large values of $B_{\perp}$, beyond roughly $3.4T$ for the parameters chosen, the ordering is $E_{{K^{\prime}\uparrow}}<E_{{K\downarrow}}<E_{{K\uparrow}}$.
@@
Figure 1: HF energies of the occupied levels for bulk $\nu=1$ as a function of ${B_{\perp}}/B_{\perp}^{0}$ with $V_{ex}=0$$E_{Z}^{0}=0.5$$E_{V}=5$$u_{z}^{0}=4$ and $u_{xy}^{0}=-3$ i.e. the case when $E_{V}>~{}(u_{z}^{0}-|u_{xy}^{0}|)>~{}E_{Z}^{0}$. For ${B_{\perp}}/B_{\perp}^{0}\simeq 1$, the ordering of the occupied levels is given by $E_{K\downarrow}<E_{K\uparrow}<E_{K^{\prime}\uparrow}$. As we increase ${B_{\perp}}$, the energy of the state ${K^{\prime}\uparrow}$ decreases linearly while the other energies increase, thus changing the order of the occupied levels.
@@
As a second illustrative example, let us consider a case where, at $E_{Z}=E_{V}=0$, the system would have been in the BO state in the physical region of anisotropic couplings, that is, $u_{z}<|u_{xy}|$. We will consider a sceniario similar to the one in the previous paragraph, with $E_{V}>(|u_{xy}^{0}|-u_{z}^{0})>E_{Z}^{0}$. Now there are only two different types of orderings of single-particle energy levels, as seen in Fig.  (2). For $B_{\perp}<E_{V}/u_{z}^{0}$ (roughly $2.5T$ for the parameters chosen) the ordering is $E_{K\downarrow}<E_{K\uparrow}<E_{K^{\prime}\uparrow}$. For larger $B_{\perp}$ the ordering switches to $E_{K\downarrow}<E_{K^{\prime}\uparrow}<E_{K\uparrow}$.
@@
Figure 2: HF energies of the occupied levels for bulk $\nu=1$ as a function of ${B_{\perp}}/B_{\perp}^{0}$ with $V_{ex}=0$$E_{Z}^{0}=0.5$$E_{V}=5$$u_{z}^{0}=2$ and $u_{xy}^{0}=-3$ i.e. the case when $E_{V}>(|u_{xy}^{0}|-u_{z}^{0})>E_{Z}^{0}$. For ${B_{\perp}}/B_{\perp}^{0}\simeq 1$, the ordering of the occupied levels are again $E_{K\downarrow}<E_{K\uparrow}<E_{K^{\prime}\uparrow}$, but as we increase ${B_{\perp}}$, the energies of all the occupied levels increase. We find that the ordering of the occupied levels changes only once, in contrast to Fig. (1) where it changes twice.
@@
From these examples it is clear that for fixed $E_{V}$ and purely perpendicular field, the ratio of $u_{z}$ and $u_{xy}$ (that is, whether $\frac{u_{z}}{|u_{xy}|}>1$ or $\frac{u_{z}}{|u_{xy}|}<1$) as well as the magnitude of $B_{\perp}$ control the energetic ordering of filled levels at $\nu=1$. While any static ground state average will not be affected by the ordering of the filled levels, we will show that this ordering has drastic effects on the junctions, and thence on the magnon transmission.